Graph the line with slope 3/4 passing through the point (3,-4) .
Answer by LinnW(1048)

Use y = mx + b as a template
m = slope = 3/4. Use the point (3,-4) meaning x=3 and y= -4
Substitute into y = mx + b
-4 = (3/4)(3) + b
-4 = 9/4 + b
add -9/4 to each side
-4 -9/4 = b
-16/4 -9/4 = b
-25/4 = b
So using y = mx + b again
we have y = (3/4)x -25/4